OpenStudy (anonymous):

Some help with simplifying. (5/a^2) - (4/(a*b)) + (7/b^2)

OpenStudy (anonymous):

\[\left( \frac{ 5 }{ a^2 } - \frac{ 4 }{ ab } + \frac{ 7 }{ b^2 } \right)\]

OpenStudy (anonymous):

I just need help starting it. I don't want you to give me the answer just like that

OpenStudy (jdoe0001):

I dont' see much factoring there btw

OpenStudy (jdoe0001):

5 and 7 are primes 4 has two factors, neither one is either 5 or 7 and the denominators can only be paired with another, but not all three

OpenStudy (anonymous):

Hmm well it appears the question was simplifying, any idea how I can simplify it?

OpenStudy (skullpatrol):

You have an expression with three terms. To simplify it you want to make into an expression with one term.

OpenStudy (skullpatrol):

Hint: common denominator

OpenStudy (anonymous):

Thanks, I think i got it :)

OpenStudy (skullpatrol):

Thanks for asking :-)